evans12124:

The high school soccer team had a Saturday car wash. They charged $3.50 for large cars and $3.00 for small cars. They had a successful day, waged a total of 124 cars and made $403.50. How many large cars and small cars did they wash?

Vocaloid:

have to set up a system, one that represents the number of cars and the other that represents the total revenue in general, for selling any number (z) items for ("b") price, we multiply price times items to get bz so let's apply this logic to the problem total number of cars = L + S = 124 (given) revenue from small cars = S * 3.00 revenue from big cars = L * 3.50 setting up a system gives us L + S = 124 3S + 3.5L = 403.50 solve for L and S. I would recommend multiplying the first equation by 3, then eliminating L through subtraction, solving for S, then going back to solve for L